what are the formulae regarding refraction of lens & mirrors

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

In mirrors, there is no refraction only reflection:(1/u)+(1/v)=1/f, f=R/2

lens makers formula:(1/f)=(n-1)((1/r1)-(1/r2))

(1/u) +(1/v)=1/f